OpenStudy (anonymous):

How many of 1, 2, 3, ...,1000 can be written ad the difference of the squares of two non-negative integers?

OpenStudy (anonymous):

This problem has bugged me few months ago, here is the M.SE thread: http://math.stackexchange.com/questions/56649/

OpenStudy (anonymous):

750

OpenStudy (anonymous):

Hey, satt, do you know any closed form?

OpenStudy (anonymous):

but i cheated. however, if you want a nice worked out solution...

OpenStudy (anonymous):

You used a program?

OpenStudy (anonymous):

no i just looked it up. the solution is comprehensible though

OpenStudy (anonymous):

Yep, check out André's answer in that url I posted.

OpenStudy (anonymous):

here is a nice brief explanation http://www.artofproblemsolving.com/Wiki/index.php/1997_AIME_Problems/Problem_1

OpenStudy (anonymous):

seems much clearer than the link at stackexchange

OpenStudy (anonymous):

For a unknown reason, my browser is failing to load the latex/mathjax in that url you posted.

OpenStudy (anonymous):

ok then i can paraphrase if you like because it is short

OpenStudy (anonymous):

That would be great :-)

OpenStudy (anonymous):

if \[x=(a-b)(a+b)\] then \[a+b,a-b\] have the same parity. so \[x\neq 4n+2=2(2n+1)\] because one factor is even and one is odd

OpenStudy (anonymous):

that leaves odd n, and for any \[x=2n+1\] you have \[x=2n+1=(n+1)^2-n^2\] and for \[x=4n\] \[x=4n=(n+1)^2-(n-1)^2\]

OpenStudy (anonymous):

I don't get it, how does this helps in counting?

OpenStudy (anonymous):

i guess the point is, that if x is odd, we have \[2n+1=(n+1)^2-n^2\] and if x is divisible by 4 we have \[4n = (n+1)^2-(n-1)^2\] and if x is divisible by 2 and not 4, then it is not possible

OpenStudy (anonymous):

so that counts out 1/4 of the numbers, leaving 750

OpenStudy (anonymous):

this is like trying to find a pythagorean triple with one side given. say the side is 7, then you write \[7=m^2-n^2=(m+n)(m-n)=7\times 1\] making \[m=4, n=3\] and the triple is \[7,24,25\] or \[m^2-m^2,2mn,m^2+n^2\]

OpenStudy (anonymous):

if the side is even but not divisible by 4, you cannot find a "primitive" triple, but you can if it is divisible by 4